If z 1=-1, z 2 = i then find Arg (z1/z2)
.

Grade:upto college level

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Arg(z1/z2)=Arg(-1/i)=Arg(i)=\pi /2
Thanks and REgards
Shaik Aasif
askIITians faculty
Indhraja
4 years ago
W.K.T Arg(-1)= π and Arg i =π÷2
Then Arg(z1/z2)=Argz1-Argz2
=π-(π÷2)=π÷2
 
Thanks and regards,
Intermediate mathematics faculty,